OpenStudy (anonymous):

write a conic equation in standard form x^2-9y^2+12x-36y+9=0

OpenStudy (anonymous):

[{(x+6)^2}/(-9)] + (y+2)^2 =1

OpenStudy (anonymous):

The goal is to complete the square for both our \(x\) and \(y\) terms.$$x^2-9y^2+12x-36y+9=0\\x^2+12x-9y^2-36y=-9\\(x^2+12x+36)-9(y^2+4y+4)=-9+36-36\\(x+6)^2-9(y+2)^2=-9\\(y+2)^2-(x+6)^2=1$$

OpenStudy (anonymous):

woops... \(-\frac{(x+6)^2}9\)

OpenStudy (anonymous):

so which is the answer ?

OpenStudy (anonymous):

...

OpenStudy (anonymous):

????